Abigail is 28 years younger than her mother Bethany. the sum of their ages is 66 . how old is each now
Answer by harpazo(655)

Let A = Abigail
Let B = Bethany

A = B - 28
A + B = 66

B - 28 + B = 66
2B - 28 = 66
2B = 66 + 28
2B = 94
B = 94/2
B = 47.
Bethany is 47.

How old is Abigail?

Think for a minute.
If I subtract 47 from the total of their ages, I will know Abigail's age, right?

So, Abigail is 66 - 47 or 19.
Abigail is 19 and Bethany is 47.